OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)? y = -2x + 1 y = 2x + 1 y = 1/2 x - 1 y = -1/2 x + 1 correct?

OpenStudy (mathstudent55):

Can you find the slope of the line using the two given pioins?

OpenStudy (anonymous):

-6/-3

OpenStudy (solomonzelman):

yes, simplify this slope....

OpenStudy (mathstudent55):

Now reduce the fraction. You can divide both the numerator and denominator by -2.

OpenStudy (anonymous):

1/-2

OpenStudy (solomonzelman):

mathstudent, you meant -3.

OpenStudy (mathstudent55):

No. First, negative divided by negative equals positive, so -6/-3 = 6/3. What is 6/3 = ?

OpenStudy (mathstudent55):

Yes, correct, thanks.

OpenStudy (anonymous):

is that it?

OpenStudy (mathstudent55):

What is 6 divided by 3 ?

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

\(slope = m = \dfrac{-3 - 3}{-2 - 1} = \dfrac{-6}{-3} = \dfrac{6}{3} = 2\)

OpenStudy (mathstudent55):

Correct. the slope is 2. The slope-intercept form of the equation of a line is y = mx + b All your choices are in this form. Ok?

OpenStudy (mathstudent55):

m, the number that multiplies x, in the slope-intercept form is the slope of the line. You know the line has a slope of 2. Which is the only choice that has a slope of 2?

OpenStudy (anonymous):

B

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

Thanks. can you help me with another equation?